During an adiabatic process the pressure of the gas is found to be proportional to fourth power of temperature. The ideal gas would be A. H2

B. He C. CH2 d micture of H2 and He




1
Expert's answer
2021-09-23T08:30:59-0400

"\alpha, \beta, \gamma" below are constants.


Ideal gas law:


"PV=nRT, \\\space\\

T=\frac{PV}{nR}=\beta PV."

According to the condition:


"P=\alpha T^4,"


and since, as we established from Ideal gas law, the temperature is proportional to the product of pressure and time, we can write


"T^4=(\beta PV)^4=
\frac{P}{\alpha}.\\\space\\
P^3V^4=\frac{1}{\alpha \beta}."

Raise both sides of the equation to the power of 1/3:


"PV^{4/3}=\frac{1}{(\alpha\beta)^3}."


An adiabatic process is described by


"PV^\gamma=\text {const}."


Since the right part of the previous equation is a constant, comparing it with the adiabatic equation, we see that


"\gamma=\frac43."

Express the adiabatic constant in terms of degree of freedom f:

"\gamma=1+\frac{2}{f},\\\space\\
\frac43=1+\frac 2f,\\\space\\
f=6"

degrees of freedom, which corresponds to

The answer is CH2